OpenStudy (anonymous):

Given that f(x) = −x + 4 and g(x) = −2x − 3, solve for f(g(x)) when x = 2. 1. -14 2.-7 3.7 4.11

Vocaloid (vocaloid):

first step is to solve g(x) when x = 2 then take the result and plug it into f(x)

OpenStudy (anonymous):

Do I divide 2 to -2x and -3? @Vocaloid

Vocaloid (vocaloid):

... huh? let's take it one step at a time: what is g(x) when x = 2?

OpenStudy (anonymous):

okay so g(2)=-2x-3 what do I do next with this funtion? @Vocaloid

Vocaloid (vocaloid):

plug in 2 for x and tell me what you get when you solve it

OpenStudy (anonymous):

I'm a bit unsure if it's right but negative 7

Vocaloid (vocaloid):

yup! g(2) = -7 alright, so now let's plug in -7 for x using f(x)

OpenStudy (anonymous):

f(-7)=-4(-7)=9 I'm I wrong at some part? @Vocaloid

Vocaloid (vocaloid):

uh, not quite f(x) = -x+4, remember?

Vocaloid (vocaloid):

if x = -7, what is f(x)?

OpenStudy (anonymous):

oh that was my fault I plugged it in for the wrong problem @Vocaloid

OpenStudy (anonymous):

instead of the x it would be F(-7) Am I right? @Vocaloid

Vocaloid (vocaloid):

yup, so what is our answer? :)

OpenStudy (anonymous):

F(-7)=-7+4 F(-7)=-3 what am I going to use F(-7) for? @Vocaloid

Vocaloid (vocaloid):

not quite, check your math again f(-7) = -(-7) + 4 = ?

OpenStudy (anonymous):

F(-7)=-7+4 F(-7)=-3 F(g(2))=-7 right? :) @vocaloid

Vocaloid (vocaloid):

uh... not quite... what is -(-7)?

Vocaloid (vocaloid):

hint: -(-a) = a

OpenStudy (anonymous):

the answer for g(x) function

Vocaloid (vocaloid):

no, can you please answer my question? -(-7) = ? hint: -(-a) = a

OpenStudy (anonymous):

I do not know

OpenStudy (anonymous):

I'm a bit slow right now sorry

Vocaloid (vocaloid):

well, ok, we know that f(-7) = -(-7) + 4, right? -(-7) = 7, so f(x) = 7 + 4 = 11 11 is our answer

Vocaloid (vocaloid):

full explanation: the question asked for f(g(x)) when x = 2 that means, plug in 2 for x to find g(2) g(2) = -7, so we plug in -7 for x in f(x) f(-7) = 11, meaning f(g(2)) = 11

OpenStudy (anonymous):

but how did you get that negative in front of the perentheses

Vocaloid (vocaloid):

f(x) = -x+4 x = -7 f(x) = -(-7) + 4

OpenStudy (anonymous):

oh so the negative of the other x was still be in play

Vocaloid (vocaloid):

yup!

OpenStudy (anonymous):

that's what i was missing hahah:) i might be a dumbarse but thanks a lot you're a really good helper :)

OpenStudy (anonymous):

@Vocaloid